8. The moon has a mass of 7.32 × 1022kg and a radius of 1609.4 km. Calculate the value of "𝑔" at the surface of the moon. 


1
Expert's answer
2021-09-23T08:33:04-0400

By the definition of the acceleration due to gravity, we have:


"g=\dfrac{GM}{R^2}=\dfrac{6.67\cdot10^{-11}\ \dfrac{N\cdot m^2}{kg^2}\cdot7.32\cdot10^{22}\ kg}{(1.6094\cdot10^6\ m)^2}=1.88\ \dfrac{m}{s^2}."